OpenStudy (anonymous):

Is this a parabola, hyperbola, ellipse, or circle? a.) y^2 + 2x^2 - 10 = 0

OpenStudy (anonymous):

please help!!!

zepdrix (zepdrix):

Do you remember what the `standard form` for each parabola, hyperbola, ellipse and circle look like? :) We'll want to take this function and put it into one of those standard forms so we can easily identify which one it is.

zepdrix (zepdrix):

\[\large y^2+2x^2-10 = 0\]Adding 10 to both sides gives us,\[\large y^2+2x^2=10\]Divide both sides by 2,\[\large \frac{y^2}{2}+x^2=5\] Hmm let's write it like this, and see if it looks familiar.\[\large \frac{y^2}{2}+\frac{x^2}{1}=5\]

OpenStudy (anonymous):

is it an ellipse?

zepdrix (zepdrix):

Ah yes it is! I should have gone one more step to make the right side 1, my mistake.\[\large \frac{y^2}{10}+\frac{x^2}{5}=1\] But yes, very good, ellipse. :)

OpenStudy (anonymous):

thank you